Find the general solution: xy – 4y = 0

please show all work.
Answer #1

solution xyl _ uyl=0 let y=x W y= mom-1) Y = m/m) 2) M-2 Put these in above differential equation a(n(n-1) 8 ) - 4(m 2-1)mesom so m/M-5) = 0 m=0 or m=5 y = 6,20 +6245 Y(X) = 41+4725Alternative solution is below Here aal, 6=-4, C = 0 Solution Xyll_uyl=0 let 2= logx Now comparing with Standard equation for Cauchy fulers method il axy0=(5-46 Jul sou ll 0 = W M) = 0,02 + کے ) OR Y(*= 6, +6,25